OpenStudy (anonymous):

determine wether the givenfunction is even odd neither f(x)= x^3-4x

OpenStudy (anonymous):

A function is even if \(f(-x)=f(x)\), odd if \(f(-x)=-f(x)\), and neither if none of these are true. If \(f(x)=x^3-4x\), then \(f(-x)=(-x)^3-4(-x)\). How does the sign compare to \(f(x)\)?

OpenStudy (anonymous):

you would have opposite signs for both x and 4 so odd

OpenStudy (anonymous):

That's right.